OpenStudy (anonymous):

i need to graph this equation but I need to solve the equation to get the points in order to do the graph y+6=0

OpenStudy (amistre64):

-6 from each side

OpenStudy (amistre64):

y+6 = 0 -6 -6 --------- y = -6

OpenStudy (anonymous):

ok, thats what I got too. so my first points would be (0,-6)?

OpenStudy (anonymous):

candrie: add 6 to each side of the equation to balance your equation. Your equation should then read y=-6. Your graph will be a straight line at the point y=-6 and will run parallel to the x-axis for infinity.

OpenStudy (anonymous):

so I will just have a vertical line?

OpenStudy (anonymous):

horizontal. should run parallel to the x-axis.

OpenStudy (anonymous):

sorry , i meant that. I was looking at the word "VERTICAL" on another problem and typed it...lol thanks